OpenStudy (anonymous):

A bag contains 5 red marbles, 6 green marbles, and 7 yellow marbles. if three marbles are randomly removed from the bag, what is the probability that they would all be green

OpenStudy (anonymous):

probability first one is green is \(\frac{6}{18}\) probability second one is green given the first one is green is \(\frac{5}{17}\) and finally the probabilty the third one is green given that the first two are green is \(\frac{4}{16}\) multiply these together to get our answer you should probably reduce first

OpenStudy (anonymous):

so 1/3 x 5/17 x 1/4